# Learning to Solve and Verify: A Self-Play Framework for Code and Test Generation

### v3

@@ 3.1 Problem Formulation
[…]
@@ 4 Experiments
Table 1: Evaluation results over training iterations for our Solver-Verifier ( Sol-Ver) method. $\Delta$(%) means the relative percentage change from Baseline to Iter3${}_{\text{+DPO}}$. Passmeans average pass rate, Accis accuracy, and FPis false positive rate. Sol-Ver provides large gains over the baseline on both code generation (solver) and unit test generation (verifier) tasks, which increase across iterations.
Baseline Iter1${}_{\text{+SFT}}$ Iter1${}_{\text{+DPO}}$ Iter2${}_{\text{+SFT}}$ Iter2${}_{\text{+DPO}}$ Iter3${}_{\text{+SFT}}$ Iter3${}_{\text{+DPO}}$ $\Delta$(%)
@@ 4.4 Evaluating Agreement between Iterations
Dataset Acc%${}_{\text{Iter1}}$ Acc%${}_{\text{Iter2}}$ Agreemnt Acc%${}_{\text{Ens.}}$
@@ 6 Conclusion
Acknowledgment
We thank Ilia Kulikov for assisting with the coding and answering questions regarding the usage of fairseq2.
@@ Appendix A Prompt Template
Prompt for Generating Problem Description [Code Snippet] for (first_p, second_p) in zip_longest(diag1, diag2): assert first_p[0] == pytest.approx(second_p[0]) assert first_p[1] == pytest.approx(second_p[1]) [Template] Polycarp is reading a book consisting of $n$ pages numbered from $1$ to $n$. Every time he finishes the page with the number divisible by $m$, he writes down the last digit of this page number. For example, if $n=15$ and $m=5$, pages divisible by $m$ are $5,10,15$. Their last digits are $5,0,5$ correspondingly, their sum is $10$. Your task is to calculate the sum of all digits Polycarp has written down. You have to answer $q$ independent queries. [Instruction] Please gain inspiration from the previous random code snippet and template to create a high-quality python programming problem. Rules: - Never mention the “code snippet”. - Don’t write the solution. - Do not specify constraints nor example inputs/outputs. - The inspiration is just an inspiration. You can deviate from it. - The solution of the problem should be only one function, not an entire program. - The problem should be self-contained.

### v4

@@ 3.1 Problem Formulation
For the verifier perspective, similarly, for each problem $P$, and each chosen tuple $(P,C,\mathbf{T}^{+}):y=1$, find a rejected tuple $(P,C,\mathbf{T}^{-}):y=0$ [^2] This can be achieved by selecting any expected output that is not equal to the chosen one in the sampling space for $T$.. The verifier-related DPO loss is then:
@@ 4 Experiments
[…]
Baseline $_{\text{+SFT}} Iter1$$_{\text{+DPO}}$ $_{\text{+SFT}} Iter2$$_{\text{+DPO}}$ $_{\text{+SFT}} Iter3$$_{\text{+DPO}}$ $\Delta$(%)
@@ 4.4 Evaluating Agreement between Iterations
[…]
@@ Appendix A Prompt Template
Prompt for Generating Problem Description [Code Snippet] for (first_p, second_p) in zip_longest(diag1, diag2): assert first_p[0] == pytest.approx(second_p[0]) assert first_p[1] == pytest.approx(second_p[1]) [Template] Polycarp is reading a book consisting of $n$ pages numbered from 11 to $n$. Every time he finishes the page with the number divisible by $m$, he writes down the last digit of this page number. For example, if $n=15$ and $m=5$, pages divisible by $m$ are $5,10,15$. Their last digits are $5,0,5$ correspondingly, their sum is 1010. Your task is to calculate the sum of all digits Polycarp has written down. You have to answer $q$ independent queries. [Instruction] Please gain inspiration from the previous random code snippet and template to create a high-quality python programming problem. Rules: - Never mention the “code snippet”. - Don’t write the solution. - Do not specify constraints nor example inputs/outputs. - The inspiration is just an inspiration. You can deviate from it. - The solution of the problem should be only one function, not an entire program. - The problem should be self-contained.